# A New Approach to Solving SMAC Task: Generating Decision Tree Code from Large Language Models

### v1

@@ Abstract
StarCraft Multi-Agent Challenge (SMAC) is one of the most commonly used experimental environments in multi-agent reinforcement learning (MARL), where the specific task is to control a set number of allied units to defeat enemy forces. Traditional MARL algorithms often require interacting with the environment for up to 1 million steps to train a model, and the resulting policies are typically non-interpretable with weak transferability. In this paper, we propose a novel approach to solving SMAC tasks called LLM-SMAC. In our framework, agents leverage large language models (LLMs) to generate decision tree code by providing task descriptions. The model is further self-reflection using feedback from the rewards provided by the environment. We conduct experiments in the SMAC and demonstrate that our method can produce high-quality, interpretable decision trees with minimal environmental exploration. Moreover, these models exhibit strong transferability, successfully applying to similar SMAC environments without modification. We believe this approach offers a new direction for solving decision-making tasks in the future.
$\dagger$\$$\dagger$\$footnotetext: This work was done when Yue Deng was an intern at Polixir
@@ 1 Introduction
The StarCraft Multi-Agent Challenge (SMAC)(Samvelyan et al., 2019) has emerged as a widely adopted benchmark for the evaluation of multi-agent reinforcement learning (MARL) algorithms. Built upon the real-time strategy game StarCraft II, SMAC provides a highly challenging environment in which each agent controls a distinct unit on the battlefield. Agents must operate in partially observable and dynamic environments, where effective coordination is crucial to achieving common objectives. With access only to local information, agents are required to collaborate while contending with adversarial units governed by built-in AI opponents. The environment offers a diverse set of tasks, ranging from small-scale skirmishes to large-scale, complex battles, making it an ideal testbed for assessing agent cooperation, coordination, and competitive behavior in MARL settings. Consequently, SMAC has become a standard benchmark for evaluating the performance of cutting-edge algorithms, including VDN(Sunehag et al., 2017), QMIX(Rashid et al., 2018), QPLEX(Wang et al., 2021a), and MAPPO(Yu et al., 2022), which have shown their effectiveness in this domain.
@@
  • Lack of model interpretability: Since deep learning relies on complex neural networks, MARL models are often considered black boxes, making it difficult to explain the decision-making process. In certain applications, such as autonomous driving or healthcare, this lack of interpretability can raise concerns about safety or regulatory compliance.
@@
Traditional rule-based decision tree models can indeed overcome the aforementioned issues, but building such models requires a significant amount of human expertise and prior knowledge, which leads to high labor costs. For a long time, there has been no effective method to integrate the strengths of both approaches.
With the advent of Large Language Models (LLMs)(Ouyang et al., 2022; OpenAI, 2023; Touvron et al., 2023; Bai et al., 2023; Du et al., 2022), especially code-oriented LLMs(Guo et al., 2024; Hui et al., 2024), the ability of models to generate code has become increasingly robust, providing a novel approach for their integration. During pre-training, LLMs learn from vast amounts of human experience and knowledge, particularly in the domain of coding, laying a strong foundation for models to generate code that can solve complex tasks. Agents can take task descriptions as input, generate corresponding decision tree code, and iteratively improve the decision tree based on task feedback. In this paradigm, not only is the challenge of lacking a reward function during LLM training addressed, but it also overcomes the issue of the non-interpretability of policies generated by reinforcement learning.
Taking the SMAC task, we propose a novel approach called LLM-SMAC. We are excited to report that LLM-SMAC, with minimal environmental exploration, produces a high-quality white-box decision tree model on the majority of SMAC maps. Moreover, this model exhibits strong transferability, as it can be applied to similar SMAC environments without any modifications. Through visualization, we observe that the agents trained with LLM-SMAC acquire the micro-skills required for SMAC, such as kite and focus fire. Moreover, these agents prioritize macro tactics and are less prone to exploiting specific loopholes in the game environment to achieve victory, resulting in higher strategy robustness.
We believe this presents a new perspective for addressing decision-making tasks. At the same time, we recognize that this is only the first step in this direction. Looking ahead, we plan to fine-tune the LLM using reward functions to further enhance the model’s capability in generating decision tree code.
@@ 2.2 LLM Agents in Complex Game Environments
Video game applications have been particularly diverse and impressive. The MineDojo environment (Fan et al., 2022) facilitated projects like GITM (Zhu et al., 2023) and Voyager (Wang et al., 2023) in Minecraft, demonstrating LLM agents’ ability to navigate and perform tasks in complex 3D environments. PokéLLMon (Hu et al., 2024) showcased their effectiveness in turn-based tactical games, emphasizing the importance of external knowledge retrieval and iterative policy refinement.
The Cradle framework (Tan et al., 2024) introduced a novel approach allowing LLM agents to interact with various software and games through a unified interface of screenshots and keyboard/mouse inputs. This demonstrated the potential for general-purpose game-playing agents across multiple commercial video games and software applications. CivRealm (Qi et al., 2024) presented a complex, imperfect-information general-sum game environment, challenging both learning and reasoning capabilities of AI agents.
@@ 3 Method
This section introduces our LLM-SMAC framework for solving SMAC tasks via decision tree codes generated by LLMs. By leveraging the ability of reasoning and planning of LLM and the code generation ability of coder LLM, our framework successfully solves most of the SMAC micro-management tasks. Figure 1 illustrates the overview of our framework and we will now proceed to elaborate on its details.
Figure 1: The overall architecture of our proposed LLM-SMAC framework. The framework takes the information of units, maps, and tasks as the environment prompts. Then the planner generates a strategy conditioned on the environment and the coder generates Python scripts following the strategy. The scripts are tested on the SMAC environment and the results with the strategy and the coder are fed into the critic to analyze the refinement suggestion. Finally, the planner and the coder generate new strategies and codes in closed loops.
@@ 3.1 Architecture
[…]
@@
To solve an SMAC task, unit and map information are analyzed and presented as an environment prompt for the planning LLM. Additionally, the current learned skill set and historical strategy traces are provided to the planning LLM to help generate a strategy skeleton. The planner then suggests skill options, describes each skill, and defines the conditions for using them. This decision-tree representation serves as the input for the LLM coder, which generates the corresponding Python scripts.
Given the environment prompt and strategy skeleton, the coder model generates a Python function that adheres to the requirements of python-sc2. Due to LLM hallucinations and the existence of multiple ways to implement a strategy skeleton, the coder iteratively generates and tests the code on SMAC maps. Based on the correctness and performance of the code, the SMAC simulation process provides stack traces of bugs or key results, such as win rates, scores, and damage statistics.
The critic module evaluates the generated code. It takes the rollout results and the script as input, with the critic LLM analyzing the reasons for high or low performance or identifying specific bugs. In addition, the critic provides suggestions for refining the strategy, as well as code corrections when bugs are present. The critic LLM also offers final recommendations on whether to revise the planned strategy or improve the current implementation, guiding the workflow for the next round, either to the planner or the coder.
3.2 Strategy planning
Figure 2: The planner generates a strategy based on the historical strategies and their scores. Meanwhile, the information on maps, units, and descriptions of tasks are also fed into the planner for more precise planning.
@@ 3.2 Strategy planning
The planner provides the coder with an overall strategy and detailed descriptions of the corresponding skills. Due to the nature of open-loop control, the planner can only offer general, coarse-grained strategy suggestions without detailed information about the agent units or the maps. These coarse-grained strategies focus more on broad strategic considerations rather than micro-management. Therefore, the planner must refine these suggestions based on unit and map information. In our framework, unit details are pre-scraped from Liquipedia, where information about unit characteristics and usage can be automatically retrieved and incorporated into the environment prompt. Meanwhile, map information, including size, available operational areas, and terrain features, is also summarized and added to the prompt.
@@
$ST=(s_{0},s_{1},\dots,s_{n})=LLM_{planner}(I_{unit}+I_{map}+h(s_{i},r_{i})),$ (1)
where the output strategy $ST$, composed of a sequence of skills $s_{i}$, is generated by the LLM planner, which takes into account unit information $I_{unit}$, map information $I_{map}$, and historical data. The historical data includes each skill $s_{i}$ in the current skill pool and the associated score $r_{i}$ for each skill.
3.3 Decision-tree Generation
Given the strategy generated by the planner, the coder module is tasked with producing the corresponding Python code using the python-sc2 package for testing. However, the performance of the generated code depends on several factors: the strategy provided, the interpretability of the LLM itself, and the quality of the training data. To enhance the strategy’s performance and facilitate the translation of the strategy into code, the planner must not only provide the skills but also specify the conditions for invoking each skill through detailed prompts:
@@ 3.3 Decision-tree Generation
The generated code is then evaluated using StarCraft II simulations running in parallel. The average win rate and additional metrics such as scores, damage dealt, damage taken on health, and damage absorbed by shields are calculated over 10 simulation runs. If any exceptions occur during testing, the stack traces of these exceptions are provided to the critic LLM for analysis, and the win rate is set to 0.
@@ 3.4 Perfomance Analyzer
If no exceptions occur during the rollout, the test results are provided to the critic model, which analyzes the reasons behind the current performance, identifying both the factors contributing to high performance and lessons from any underperformance. Even when exceptions are present, valuable insights can still be gained. Ultimately, the critic provides explanations, potential improvements, and recommendations on whether to adjust the strategy or refine the current approach ($LLM_{critic}(C^{\prime},G)$) to guide the next steps in the pipeline.
@@
$r_{i}=\frac{(r_{i}\times\Gamma_{i}+G)}{\Gamma_{i}+1}$ (4)

### v2

# SMAC-R1: The Emergence of Intelligence in Decision-Making Tasks

@@ Abstract
StarCraft Multi-Agent Challenge (SMAC) has been one of the most commonly used experimental environments in multi-agent reinforcement learning (MARL), where the specific task is to control a set number of allied units to defeat enemy forces. Traditional MARL algorithms often require interacting with the environment for millions of steps to train a parametric model, of which the resulting policies are typically non-interpretable with weak transferability. In this paper, we introduce SMAC-R1 which is based on the Qwen2.5-7B-Base LLM distilled from DeepSeek-Coder-v2.5-236B. Similar to online reinforcement learning after behavior cloning in offline learning process, in our pipeline, agents leverage the DeepSeek LLM to generate decision tree code by providing task descriptions, and the agents are further self-reflected using feedback from the rewards provided by the environment. Based on that, we augment the generated scripts to fine-tune a small LLM, Qwen2.5-7B-Base, to distill the decision-making ability via Supervised Fine-Tuning (SFT) and enhance the script generation ability by the Group Relative Policy Optimization (GRPO) algorithm. We conduct experiments in the original 23 SMAC tasks and 10 newly-designed tasks to demonstrate that our method can produce high-quality, interpretable decision trees with minimal environmental exploration. Moreover, these scripts exhibit strong transferability, successfully applying to homogeneous SMAC environments without modification. We believe this approach offers a new direction for solving decision-making tasks and domain-specific LLM training pipelines in the future.
@@ 1 Introduction
The StarCraft Multi-Agent Challenge (SMAC)(Samvelyan et al., 2019) has emerged as a widely adopted benchmark for the evaluation of multi-agent reinforcement learning (MARL) algorithms. Built upon the real-time strategy game StarCraft II, SMAC provides a highly challenging environment in which each agent controls a distinct unit on the battlefield. Agents must operate in partially observable and dynamic environments where effective coordination is crucial to achieving common objectives. With access only to local information, agents are required to collaborate while contending with adversarial units governed by built-in AI opponents. The environment offers a diverse set of tasks, ranging from small-scale skirmishes to large-scale, complex battles, making it an ideal testbed for assessing agent cooperation, coordination, and competitive behavior in MARL settings. Consequently, SMAC has become a standard benchmark for evaluating the performance of cutting-edge algorithms, including VDN(Sunehag et al., 2017), QMIX(Rashid et al., 2018), QPLEX(Wang et al., 2021a), and MAPPO(Yu et al., 2022), which have shown their effectiveness in this domain.
@@
  • Lack of model interpretability: Since deep learning relies on complex neural networks, MARL models are often considered black boxes, making it difficult to explain the decision-making process. In certain applications, such as autonomous driving or healthcare, this lack of interpretability can raise concerns about safety or regulatory compliance. In contrast, humans may benefit from white-box reasoning such as human-in-the-loop and bad case correction scenarios when the decision-making process is completely exposed to humans.
@@
Traditional rule-based decision tree models can indeed overcome the aforementioned issues in the white-box method, but building such models requires a significant amount of human expertise and prior knowledge, which leads to high labor costs. For a long time, there has been no effective method to integrate the strengths of both approaches.
With the advent of Large Language Models (LLMs)(Ouyang et al., 2022; OpenAI, 2023; Touvron et al., 2023; Bai et al., 2023; Du et al., 2022), especially code-oriented LLMs(Guo et al., 2024; Hui et al., 2024), the ability of models to generate code has become increasingly robust, providing a novel approach for their integration. During pre-training, LLMs learn from vast amounts of human experience and knowledge, particularly in the domain of coding, laying a strong foundation for models to generate code that can solve complex tasks. Agents can take task descriptions as input, generate corresponding decision tree code, and iteratively improve the decision tree based on task feedback. This paradigm not only addresses the challenge of lacking a reward function during LLM training but also overcomes the issue of the non-interpretability of policies generated by reinforcement learning.
Meanwhile, in recent advancements in reinforcement learning from human feedback (RLHF), frameworks such as LLAMA Factory (Zheng et al., 2024), OpenRLHF (Hu et al., 2024a), and VERL (Sheng et al., 2024) have emerged as pivotal tools for aligning AI systems with human preferences. OpenRLHF provides an open-source platform for implementing RLHF, enabling researchers to fine-tune models using human feedback in a scalable and efficient manner. Similarly, the VERL framework focuses on verifiable and interpretable RLHF, ensuring that the learned policies are not only effective but also transparent and robust. Complementing these frameworks, we distill the problem-solving ability from the DeepSeek-Coder-V2.5-236B to small Qwen2.5-Coder-7B, thereby reducing computational costs while maintaining performance. Together, these approaches represent significant strides in creating AI systems that are both aligned with human values and practical for real-world deployment.
Taking the SMAC task, we propose our SMAC-R1 model based on Qwen2.5-7B-base, through the SFT and GRPO enhancement pipeline after the plan-code-critic process. We are excited to report that SMAC-R1, with minimal environmental exploration, produces a high-quality white-box decision tree model on the majority of the original 23 SMAC maps and 10 newly-designed maps. Moreover, this model exhibits strong transferability, as it can be applied to similar few-shot SMAC environments without any modifications. Through visualization, we observe that the agents provided by our SMAC-R1 model acquire the micro-skills required for SMAC, such as kite and focus fire. Moreover, these agents prioritize macro tactics and are less prone to exploiting specific loopholes in the game environment to achieve victory, resulting in higher strategy robustness. Additionally, focusing on the SMAC coding tasks, the GRPO results shows completely opposite results compared to DeepSeek models that the response body containing strategy and skills tend to be short yet precise for SMAC instead of long answers for math-solving problems. We believe this presents a new perspective for addressing decision-making tasks.
@@ 2.2 LLM Agents in Complex Game Environments
Video game applications have been particularly diverse and impressive. The MineDojo environment (Fan et al., 2022) facilitated projects like GITM (Zhu et al., 2023) and Voyager (Wang et al., 2023) in Minecraft, demonstrating LLM agents’ ability to navigate and perform tasks in complex 3D environments. PokéLLMon (Hu et al., 2024b) showcased their effectiveness in turn-based tactical games, emphasizing the importance of external knowledge retrieval and iterative policy refinement.
The Cradle framework (Tan et al., 2024) introduced a novel approach allowing LLM agents to interact with various software and games through a unified interface of screenshots and keyboard/mouse inputs. This demonstrated the potential for general-purpose game-playing agents across multiple commercial video games and software applications. CivRealm (Qi et al., 2024) presented a complex, imperfect-information general-sum game environment, challenging both the learning and reasoning capabilities of AI agents.
@@ 3 Method
This section presents the training pipeline of our proposed SMAC-R1 model designed to address SMAC tasks, including generating decision tree codes from coder LLMs, distilling decision-making ability to smaller student LLM, using generated scripts, and enhance reasoning capabilities via the GRPO algorithm. By harnessing the reasoning, planning, and code-generation capabilities of the DeepSeek LLM, our framework effectively solves the majority of SMAC micro-management tasks. Furthermore, we transfer the policy-generation ability to our SMAC-R1 model, which is based on the Qwen2.5-7B-Base, through SFT and the GRPO algorithm. The training pipeline, depicted in Figure 1 and Figure 2, encompasses script generation, script augmentation, SFT training, and GRPO-based improvement. The subsequent sections will provide a detailed explanation of these components.
@@ 3.1 Architecture
Overview
LLMs have acquired extensive knowledge about StarCraft II and learned python-sc2 and other SC2 scripts during their pretraining phase. Building upon this foundation, our framework addresses the code generation task for the python-sc2 package [^1] https://github.com/BurnySc2/python-sc2, which provides a comprehensive wrapper for StarCraft II gameplay. Unlike traditional MARL models on the pysc2 platform, the scripts from the python-sc2 package are open-loop control scripts that require the LLM to plan and anticipate game scenarios before execution. The extensive workload of planning, code generation, bug identification, and bug fixing places significant strain on a single LLM. To address this, in the decision-tree generation phase, our framework divides the process into three critical modules: the strategy planning model, the decision-tree code generation model, and the summarization critic model. Then the successfully generated scripts that can win the SMAC task are augmented to increase the diversity and used to SFT and DPO a student LLM. The SFT and DPO processes provide student LLMs standard architecture of response script, which should prevent from API mis-invocation and async function definition. Then the fine-tuned LLMs are further improved by GRPO by the actions sampled from the student itself to increase the winning rate of generated scripts.
Decision Tree Script Generation
@@
As illustrated in Figure 1, to solve a SMAC task, unit and map information are analyzed and presented as an environment prompt for the planner LLM. Additionally, the current learned skill set and historical strategy traces are provided to the planning LLM to help generate a strategy skeleton. The planner then suggests skill options, describes each skill, and defines the conditions for using them. This decision-tree representation serves as the input for the LLM coder, which generates the corresponding Python scripts. Due to LLM hallucinations and the existence of multiple ways to implement a strategy skeleton, the coder iteratively generates and tests the code on SMAC maps. Based on the correctness and performance of the code, the SMAC simulation process provides stack traces of bugs or key results, such as win rates, scores, and damage statistics. These results and the scripts serve as the input to the critic module to analyze the reasons for high or low performance or identify specific bugs. Then, the critic provides suggestions for refining the strategy, as well as code corrections when bugs are present. The critic LLM also offers final recommendations on whether to revise the planned strategy or improve the current implementation, guiding the workflow for the next round, either to the planner or the coder.
SFT and GRPO
During the distillation and fine-tuning phase, we utilize synthetically generated Python scripts alongside a coder LLM to produce novel decision tree-based scripts. A secondary coder LLM is employed to strategically rewrite these scripts, preserving their core tactical logic while altering their implementation details. The revised scripts are then evaluated in SMAC environments through rollout simulations, where they are categorized into three classes: good cases (high win rate), bad cases (low win rate), and bug cases (containing bugs identified via stack traces). To mitigate data imbalance, duplicate scripts are eliminated by comparing their abstract syntax tree (AST) representations. Our SMAC-R1 model is subsequently trained in two stages: first through supervised fine-tuning (SFT) on the successful scripts, then via Direct Preference Optimization (DPO) using both successful and suboptimal scripts to minimize API misuse risks. Then, the Group Relative Policy Optimization (GRPO) algorithm refines the model’s code generation capability, aligning it with task-specific performance metrics in SMAC scenarios.
3.2 Planner-Coder-Critic
Figure 1: The overall architecture of our proposed LLM-SMAC framework. The framework takes the information of units, maps, and tasks as the environment prompts. Then, the planner generates a strategy conditioned on the environment, and the coder generates Python scripts following the strategy. The scripts are tested on the SMAC environment, and the results with the strategy and the coder are fed into the critic to analyze the refinement suggestion. Finally, the planner and the coder generate new strategies and codes in closed loops.
Planner
@@ 3.2 Planner-Coder-Critic
The planner provides the coder with an overall strategy and detailed descriptions of the corresponding skills. Due to the nature of open-loop control, the planner can only offer general, coarse-grained strategy suggestions without detailed information about the agent units or the maps in each iteration. Therefore, the planner must refine these suggestions based on unit and map information. In our framework, unit details are pre-scraped from Liquipedia, where information about unit characteristics and usage can be automatically retrieved and incorporated into the environment prompt. Meanwhile, map information, including size, available operational areas, and terrain features, is also summarized and added to the prompt.
@@
$\small ST=(s_{0},...,s_{n})=LLM_{planner}(I_{u}+I_{m}+h(s_{i},r_{i})),$ (1)
where the output strategy $ST$, composed of a sequence of skills $s_{i}$, is generated by the LLM planner, which takes into account unit information $I_{u}$, map information $I_{m}$, and historical data. The historical data includes each skill $s_{i}$ in the current skill pool and their associated score $r_{i}$.
Coder
Given the strategy generated by the planner, the coder module is tasked with producing the corresponding Python code using the python-sc2 package for testing. However, the performance of the generated code depends on several factors: the strategy provided, the interpretability of the LLM itself, and the quality of the training data that used for training the LLM. To enhance the strategy’s performance and facilitate the translation of the strategy into code, the planner must not only provide the skills but also specify the conditions for invoking each skill through detailed prompts:
@@
The generated code is then evaluated using StarCraft II simulations. The average win rate and additional metrics such as scores, damage dealt, damage taken on health, and damage absorbed by shields are calculated over 10 simulation runs. If any exceptions occur during testing, the stack traces of these exceptions are provided to the critic LLM for analysis, and the win rate is set to 0.
@@
If no exceptions occur during the rollout, the test results are provided to the critic model, which analyzes the reasons behind the current performance, identifying both the factors contributing to high performance and lessons from any underperformance. Ultimately, the critic provides explanations, potential improvements, and recommendations on whether to adjust the strategy or refine the current approach ($LLM_{critic}(C^{\prime},G)$) to guide the next steps in the pipeline.
@@
$r_{i}={(r_{i}\times\Gamma_{i}+G)}/{(\Gamma_{i}+1)}$ (4)
@@
3.3 SMAC-R1 training
Figure 2: The overall architecture of distilling, fine-tuning, and improving process of the student LLM (SMAC-R1). The generated decision-tree scripts are augmented by a coder LLM and generate a bunch of scripts with similar strategies. The scripts are divided into good, bad, and bug cases by evaluation on SMAC. Then the SMAC-R1 is fine-tuned by SFT, DPO, and finally GRPO by the data and iteratively fine-tuned by the scripts generated by our SMAC-R1 model.
According to the planner-coder-critic framework description in the previous section, successful Python script generation depends on the success of the three modules. Consequently, the failure of any module will result in the final script generation. Meanwhile, the generating difficulty also results in the insufficient amount of script file collection for fine-tuning. Therefore, we utilize another LLM to expand the generated scripts and tune our SMAC-R1 in single-step generation format instead of multi-step reasoning.
Script data augmentation
We use the Python script generated by the coder to fine-tune our SMAC-R1 via data generation, duplicated file removal, SFT based on good cases, DPO based on good cases and bad cases and GRPO according to the SMAC reward function. To do so, we first use another LLM to iteratively imitate and rewrite the provided code, which provides a bunch of new scripts. After that, the newly generated Python scripts are tested via the SMAC environment and classified into good cases (winning rate larger than 80%), bad cases (winning rate less than 80%), and bug cases.
Duplicated file removal
Because the generated codes derive from very few successful examples, these files have a certain probability of being duplicated. We design an algorithm to evaluate and minimize redundancy within code datasets by leveraging structural normalization and similarity computation. We first do the code normalization that parses source code into its Abstract Syntax Tree (AST) representation using the Python AST module. A customized AST transformer, CodeNormalizer, standardizes variable names, function names, and class names. This transformation ensures that semantically identical but syntactically varied code fragments are normalized to a common structure. Then, the normalized ASTs are converted to string representations to calculate the similarity ratio between different code fragments by structural comparison. By comparing structural representations, the algorithm outputs a similarity percentage to judge the similarity between two script files in order to control the diversity of datasets. This approach ensures that dataset redundancy is analyzed at a structural level, providing a robust method for identifying semantically equivalent code with minimal syntactic influence.
Training Pipeline
After removing the redundancy script code, we firstly fine-tune the SMAC-R1 base model (Qwen2.5-7B-Base) by the good cases via OpenRLHF (Hu et al., 2024a) package. To mitigate the problem of API misuse, the signatures, the usage, and the effects of the functions described in the official documentation are crawled, formatted and added to the training dataset. Apart from that, we also crawl the python-sc2 scripts from the aiarena where the StarCraft community players upload their mandatory edited scripts, and provide the codes to LLM to learn other API invocation. After the SFT process, we continue to fine-tune the reference model by DPO algorithm according to the equation:
$\mathcal{L}_{DPO}=-\mathbb{E}_{(x,y_{c},y_{r})\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{c}|x)}{\pi_{ref}(y_{c}|x)}-\beta\log\frac{\pi_{\theta}(y_{r}|x)}{\pi_{ref}(y_{r}|x)}\right)\right],$ (5)
where $\mathcal{D}$ is the dataset, $\sigma$ is the sigmoid function, $\beta$ is a hyper-parameter, $y_{c}$ is the chosen response, $y_{r}$ is the rejected response, and $x$ is the input prompt.
After the DPO fine-tuning, we leverage the GRPO (Shao et al., 2024) algorithm to improve our SMAC-R1 model via reinforcement learning. According to the GRPO algorithm formation, the system prompts are treated as the question $q$, and the scripts generated are the answers $o_{i}$. Then, GRPO samples a group of outputs $\{o_{1},o_{2},...,o_{G}\}$ from the old policy $\pi_{\theta_{old}}$ in each iteration. Then our SMAC-R1, $\pi_{\theta}$, is optimized by maximizing the following objective:
$\begin{split}\mathcal{J}_{GRPO}=&\mathbb{E}_{\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{old}}(O|q)}\\ &\frac{1}{G}\sum_{i=1}^{G}(\min(\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{old}}(o_{i}|q)}A_{i},clip(\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{old}}(o_{i}|q)},1-\epsilon,1+\epsilon)A_{i})-\beta\mathbb{D}_{KL}(\pi_{\theta}||\pi_{ref}))\end{split}$ (6)
where the $\epsilon$ and $\beta$ are hyper-parameters, and $A_{i}$ is the advantage that is computed using a group of rewards $\{r_{1},r_{2},...,r_{G}\}$ corresponding to the outputs within each group:
$A_{i}=\frac{r_{i}-mean(\{r_{1},r_{2},...r_{G}\})}{std{\{r_{1},r_{2},...r_{G}\}}}$ (7)
The SMAC environment is a naturally suitable reward provider, so we define the score of the GRPO algorithm as follows:
  1. 1.
    $-1$: The valid Python script cannot be retrieved from the response content of SMAC-R1.
  2. 2.
    $-0.5$: There are bugs in the script, and the test process fails to terminate normally.
  3. 3.
    $0\sim 1$: The winning rate of the generated code in SMAC tasks.